OpenStudy (anonymous):

1. Part A.) For the f(x) = 5x + 8 and g(x) = x – 3. Calculate (f o g)(x) Part B.) f(x) = 3x + 5 and g(x) = x². Calculate (f o g)(-3) SHOW ALL WORK

Nnesha (nnesha):

Part A.) `For the f(x) = 5x + 8 and g(x) = x – 3. Calculate (f o g)(x)` fog = f composed with g of x ( fog )(x) = f(g(x))

OpenStudy (anonymous):

What f o g(x) is really saying is to find f(g(x)). For example, if you are asked for f o g(2), you would compute g(2), which would give you a number, let's pretend it is 5. Then you would stick that 5 into f and compute f(5), which is just f(g(2)) Say f(x) = 2x and g(x) = 4x + 2. f o g(x) = f(g(x)) = f(4x + 2) = (4x + 2) Try it now using that!

Nnesha (nnesha):

so substitute all x for g(x) function into f(x) equation \[\rm f(x)= 5\color{Red}{x}+8\]\[\rm g(x)=x-3\] \[\rm ( fog)(x) \rightarrow f(g(x))\]

OpenStudy (anonymous):

so whats the steps and answer, best answer ill give medals to

OpenStudy (anonymous):

I totally havve a typo in my answer, it should be 2*(4x+2)

Nnesha (nnesha):

like jtsrumps gave u an example \[\rm (f og )(2) \rightarrow f(g(2))\]

OpenStudy (anonymous):

i have idea how to do this ive been absent. where are the steps and answer

Nnesha (nnesha):

\[\rm f(x)= 5\color{Red}{x}+8\]\[\rm g(x)=x-3\] substitute x for g(x) equation which is x -3

OpenStudy (anonymous):

so whats the entire thing @Nnesha

OpenStudy (anonymous):

for A and B

Nnesha (nnesha):

okay... that's how we should replace x with x-3 \[\rm f(x)= 5\color{Red}{x}+8\] \[\rm f(\color{reD}{g(x)})= 5(\color{Red}{x-3})+8\] now distribute (x-3) by 5

OpenStudy (anonymous):

can you paste the whole thing in one post and then i can just use it plz

OpenStudy (anonymous):

or message me

Nnesha (nnesha):

no i want you to work with me

OpenStudy (anonymous):

im willing to igve you medals

Nnesha (nnesha):

i don't care about medals. i want you to learn how to solve these types of question

Nnesha (nnesha):

familiar with the distributive property ? \[\rm \color{Red}{a}(b+c)=\color{Red}{a}*b+\color{ReD}{a}*c=\color{Red}{a}b+\color{ReD}{a}c\]

OpenStudy (anonymous):

yes not functions tho

Nnesha (nnesha):

alright then distribute (x-3) by 5 let me know what you get \[\rm f(\color{reD}{g(x)})= 5(\color{Red}{x-3})+8\]

OpenStudy (anonymous):

5x-3+8?

Nnesha (nnesha):

multiply both terms in the parentheses by 5 just like in that example a times b +a times c

OpenStudy (anonymous):

5x-15+8

OpenStudy (anonymous):

5x-23

Nnesha (nnesha):

hmm how did you get 23 -15+8 = ?

OpenStudy (anonymous):

oh -7

Nnesha (nnesha):

yes that's correct that's it !

OpenStudy (anonymous):

so its -7 the answer

Nnesha (nnesha):

no

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Typical_karen 5x-15+8 \(\color{blue}{\text{End of Quote}}\) there supposed to be -7 instead 23

Nnesha (nnesha):

5x-15+8 that's what u got after distributing by 5 right that's correct and then combine like terms that's it done!

OpenStudy (anonymous):

so hoow would i put it all together

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

what do you mean ?? you can't add combine 5x with - 7 they aren't like terms

OpenStudy (anonymous):

i mean how do i put everything together to get my answer, so i can put it on paper

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\rm f(x)= 5\color{Red}{x}+8\]\[\rm g(x)=x-3\] substitute x for g(x) equation which is x -3 \(\color{blue}{\text{End of Quote}}\) \(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\rm f(\color{reD}{g(x)})= 5(\color{Red}{x-3})+8\] apply distributive property \(\color{blue}{\text{End of Quote}}\) \(\color{blue}{\text{Originally Posted by}}\) @Typical_karen 5x-15+8 \(\color{blue}{\text{End of Quote}}\) and then last step combine like terms

Nnesha (nnesha):

make sense ?

OpenStudy (anonymous):

yeah i guess

OpenStudy (hyroko):

Your answer at the ens, is 5x-7 the X doesn't go away.

Nnesha (nnesha):

alright can you start working on part B ??? that's the same question :=)) and feel free to ask any question about that

OpenStudy (preetha):

Wow, good work Nnesha!

Nnesha (nnesha):

Thanks!! :=))